# Fighting Exponentially Small Gaps by Counterdiabatic Driving

### v2

@@ V.3 Demonstration on the $3$-regular $3$-XORSAT
Finally, we also test the QBCD method on the $3$-regular $3$-XORSAT corresponding to optimization problems whose solutions can be represented as the ground state of a spin glass Hamiltonian with cubic spin interactions. In particular, the $3$-XORSAT problem can be formulated in terms of linear equations modulo $2$ with each of them containing three variables with the constraint that each variable appears in exactly three equations without repetition,
@@
Here $\underline{\underline{A}}$ is an $L\times L$ matrix with each row containing only three non-zero elements corresponding to the prescription of the problem. As the possible outcomes can only be 00 or $1$, the assigned variables can take the same values as well, $x_{i}=0,1$. The corresponding Hamiltonian is given by
$H_{P}=\sum_{c}\frac{1+J_{c}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}}{2},$ (54)
with alternating randomly as $J_{c}=\pm 1$. Following the strategy of Refs. [114, 13] we only consider such instances where the ground state is non-degenerate, allowing for a better characterization of the energy gap at the critical point. In these instances, one can drop the alternating coupling constant, and it is satisfactory to consider the Hamiltonian instead
$H_{P}=\sum_{c}\frac{1-\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}}{2},$ (55)
where the summation runs over $M=L$ equations. The corresponding quantum annealing Hamiltonian is written as
@@
exhibiting an exponentially small gap around $\lambda_{c}=1/2$, where exponentially long driving times are required to reach near adiabatic dynamics. To this end, we probe the performance of both the QBCD Hamiltonian and the first-order locally optimized variational CD ansatz. Similar to the Max Cut, we investigate the final ground state fidelity and the minimal energy gap at the critical point. The first-order CD expansion is given by
@@
with variational coefficients obtained by the solutions of the following equations (see App. H)
Within the QBCD approach, we follow the same approach as in the case of the Max Cut and determine with a mesh of solely $40$ points the minimal gap, $\Delta$ and the corresponding eigenvectors, $\lvert\mathrm{GS}\rangle$ and $\lvert\mathrm{ES}\rangle$ for the ground state and the first excited state, respectively. This method requires around two orders of magnitude fewer computational resources than simulating the annealing process. The corresponding QBCD Hamiltonian is written as
$H_{\mathrm{QBCD}}=-i\frac{\langle\mathrm{GS}\lvert\partial_{\lambda}H\rvert\mathrm{ES}\rangle}{\Delta}\left(\lvert\mathrm{ES}\rangle\langle\mathrm{GS}\rvert-\lvert\mathrm{GS}\rangle\langle\mathrm{ES}\rvert\right).$
Additionally, we also sparsified the full QBCD to the same density as that of the leading order local CD expansion, Eq. (V.3), reducing exponentially its non-locality and the computational resources of its implementation on quantum computers and in classical algorithms.
To this end, we use again the ground state gap and the final ground state fidelity as measures of adiabaticity. As shown in Fig. 16, even for short annealing times, $T=5$, the local expansion approach could only amplify the exponential decay of that ground state gap by a constant prefactor. On the other hand, the sparsified QBCD method achieves exponentially larger gap sizes with an approximately polynomial decay even for much longer annealing times, $T=10,20,80$. Its effects only vanish once adiabaticity sets in, already in the process without CD. Quite remarkably, the relative difference from the ground state gap of the bare process exhibits the same $L$-independent quadratic decay with respect to $T$ as in the case of the Max Cut, as demonstrated in Fig. 17. Similar trends are observed for the final fidelity, as shown in Fig. 18. The sparsified QBCD achieves significantly higher fidelity at short annealing times ($T=2$) than both the local expansion and the bare process, even in cases when the latter are run for much longer durations ($T=40$). At these longer time scales, the fidelity for QBCD remains close to unity, while the dynamics becomes essentially insensitive to the presence of the local CD expansion. In this regime, the local expansion together with the bare process still exhibits an exponentially decaying fidelity.
@@
Figure 17: Relative difference between the energy gaps without CD and with the first order ansatz decaying approximately as the minimal spin glass bottleneck independently of the system size. No essential gap amplification is present around the driving times, where even without CD, the fidelity departs from zero.
@@ VI Discussion
In this study, we explored how quantum adiabatic dynamics can be enhanced to overcome spin-glass bottlenecks using various approximate counterdiabatic (CD) driving schemes. Specifically, we examined the variational local expansion, its improved version known as counterdiabatic optimized local driving (COLD). We also introduced a new scheme—quantum brachistochrone counterdiabatic driving (QBCD)—which focuses directly on the bottleneck region while relying only on the approximate knowledge of an exponentially small fraction of the full CD at a single parameter value. First, we tested the local expansion approaches on an exactly solvable minimal spin glass toy model described by an Ising chain admitting a free-fermionic description, first introduced in Ref. [99]. This model serves as a valuable testbed displaying both a ground state gap, exponentially small in the system size and two degenerate ground states that differ macroscopically from each other in terms of spin flips. The solvability of the model, however, was used only to reach larger system sizes and extract the asymptotic scalings: most of our analysis can be applied without substantial modifications to NP-hard Ising Hamiltonians. To this end, we chose to employ CD terms that can be constructed efficiently on a classical computer: we resorted to two- and three-body ansatzes with coefficients fixed from the Floquet-Krylov variational principle [58, 59]. A key finding is that, while these CD ansatzes can exponentially amplify the size of the minimal gap along the adiabatic path, such gap remains exponentially small in the system size $L$, and the probability of ending up in the ground state at the end of the protocol is correspondingly very small for times shorter than $\sim e^{\alpha L}$, with $\alpha$ given in Eq. (19).
@@
Our analysis highlights key features of CD of frustrated spin systems. First, from the gap characterization in Sec. IV.1, it can be inferred that CD can amplify the gap exponentially in the diabatic regime, and thus substantially reduce the diabatic transitions. However, reaching the regime of driving times where diabatic transitions come solely from the exponentially small ground state gap, the first two CD terms provide negligible improvement (see Fig. 7) and higher order expansions are required. On the other hand, slowing down the driving schedule also reduces the overall strength of the CD term, given the $\partial_{t}\lambda$ prefactor in Eq. (21). These features prevent the construction of an effective CD term that ensures adiabatic dynamics in the presence of exponentially small gaps between frustrated energy levels.
In addition, CD techniques, including the variational formulation [58, 59, 60], aim at keeping the evolved state as close as possible to the instantaneous ground state. The simplicity of the model allows us to identify a single transition from one end to the other of the effective chain, Eq. (16). Thus, in crossing the first-order phase transition, the state needs to be modified extensively in order to adjust to the change in the ground state.
@@
As an alternative and novel approach, we have introduced the QBCD Hamiltonian in Sec. V, which shares the structure of the Hamiltonian for the time-optimal evolution in the quantum brachistochrone problem for the associated subspace of localized edge states. The QBCD Hamiltonian employs only an exponentially small fraction of the full CD at a single parameter value, focusing exclusively on the ground and first excited states at the critical point. With the help of it, the success rate of the protocol increases from a vanishing probability to almost 0.5, as indicated by the average kink number at the end of the schedule. While implementing non-local, multi-body interactions on quantum hardware remains technically challenging, the structure of QBCD makes its realization feasible with current digital or digital-analog quantum simulation platforms. This feasibility arises from two key features: only an exponentially small fraction of the full counterdiabatic correction—restricted to the ground and first excited states—is required, and an approximate knowledge of this Hamiltonian at a single parameter value is sufficient. These properties substantially reduce the complexity of the implementation, placing experimental realization of QBCD within reach. Additionally, QBCD offers a practical advantage in terms of energetic cost: the Hilbert-Schmidt norm of the Hamiltonian remains finite due to the exponentially reduced contributions from higher-order multi-body terms.
@@
In the second part of the paper, we probed the generality of our findings in two realistic spin glass models representing NP-hard problems, namely the $3$-regular Max Cut and the $3$-XORSAT problems. These models exhibit an exponentially small gap already at system sizes $L\gtrsim 10$, accessible by classical numerical simulations. Specifically, the $3$-XORSAT problem extends the validity of our results to cubic spin-glass Hamiltonians. Due to limitations in accessible system sizes, we used the final ground-state fidelity—rather than the excitation density—as a measure of adiabaticity, alongside the minimum gap. Our findings indicate that the local CD expansions (second-order expansion for Max Cut and first-order for $3$-XORSAT) display a level of effectiveness similar to that observed in the minimal spin-glass model. Only modest improvements were observed outside the adiabatic regime, in both the gap and the final fidelity. Meanwhile, for intermediate annealing times, where the fidelity departs from zero in the bare process without CD, no significant improvement was detected.
For the QBCD method, we introduced an enhanced, sparsified version that matches the density of the second-order local expansion. This sparsification significantly improves its prospects for implementations in both experimental setups and large-scale classical algorithms. Remarkably, even retaining only an exponentially small fraction of the already exponentially restricted non-local terms in the QBCD Hamiltonian yields a similar improvement in the adiabatic time scale as seen in the minimal spin-glass model.
@@
Finally, we also compared the efficiency of the COLD approach with the local expansion and the sparsified QBCD in the Max Cut problem. Although it provided a noticeable improvement over the local expansion, its performance remained consistently inferior to that of the sparsified QBCD. In fact, even at small system sizes ($L=6$), spin-glass effects already degrade COLD’s effectiveness below that of QBCD. Moreover, the COLD method proved surprisingly difficult to implement numerically—even for modest system sizes ($L\lesssim 10$). Furthermore, as the system size increased and spin-glass effects became more pronounced, the performance of COLD gradually converged to that of the second-order local expansion. These findings underscore both the limited impact of COLD and the significant computational challenges compared to the QBCD method.
@@
This combination of results suggests that the challenge of traversing exponentially small gaps in these spin-glass problems is primarily of entropic rather than energetic origin [118]. In other words, overcoming the bottleneck does not require access to strong control fields, but rather a minimal amount of information about where—and how—to steer the quantum state. This perspective further highlights the practical value of methods like QBCD, which leverage such targeted information with limited and implementable control resources.
@@
Accordingly, while exploring alternative strategies beyond a priori action minimization remains an interesting direction, our results indicate that focusing on further improving the QBCD approach may ultimately be more effective for overcoming bottlenecks. Approaches based on integrating CD ansatzes with optimal control methods—such as QAOA [63, 62, 64, 65]—optimize control fields iteratively using only the final state reached by the dynamics [77]. However, these methods often suffer from limited analytical transparency and can be hampered by complex control landscapes and barren plateaus [119, 120, 121, 122]. In contrast, QBCD already demonstrates substantial improvements with minimal and experimentally realistic information. Given these advantages, it appears more promising to build upon QBCD and approximate CD strategies—potentially enhancing them with methods such as reverse annealing [104, 105, 106], bias fields [123, 87], Lyapunov control [67, 68], or targeted control engineering near critical points.
@@ A.1 Jordan-Wigner transformation and Dirac fermions
$H[\lambda]=\sum_{j=1}^{L}\left[-\lambda J_{j}\tilde{\sigma}^{x}_{j}\tilde{\sigma}^{x}_{j+1}+(1-\lambda)\tilde{\sigma}_{j}^{z}\right]$ (62)
(tildes are used to distinguish the two sets of Pauli matrices for later convenience). This form of the Hamiltonian is convenient for a Jordan-Wigner transformation to be applied: one passes from spins $\{\tilde{\sigma}_{j}\}$ to Dirac fermions $\{c_{j}^{\phantom{\dagger}},c_{j}^{\dagger}\}$ by means of the identities
@@
Then, the fermionic vacuum |$\ket{0_{c}}$ corresponds to spins positively aligned along $x$-axis in the original basis (|$\ket{0_{c}}=\ket{\rightarrow}$), and the fully occupied fermionic state corresponds to the spins anti-aligned along $x$-axis (|$\ket{1_{c}}=\ket{\leftarrow}$).
@@
The Hamiltonian is manifestly a quadratic form in the creation/annihilation operators. In order to fix the boundary conditions, one needs to look back at the original Hamiltonian (1). In terms of the spins, the boundary conditions are periodic: $\sigma_{j+L}\equiv\sigma_{j}$. However, for the fermionic operators, it holds $c_{j+L}=+c_{j}$ or $c_{j+L}=-c_{j}$ depending on whether there is an odd or even number of fermions, respectively. Since the initial state (ground state at $\lambda=0$) is the fermionic vacuum |$\ket{0_{c}}$, the correct sector to take is the even one.

### v3

@@ V.3 Demonstration on the $3$-regular 33-XORSAT
Finally, we test the QBCD method on the $3$-regular 33-XORSAT, an optimization problem whose solutions correspond to the ground state of a spin-glass Hamiltonian with three-body interactions. In particular, the $3$-XORSAT problem can be formulated in terms of linear equations modulo $2$ with each of them containing three variables with the constraint that each variable appears in exactly three equations without repetition,
@@
Here $\underline{\underline{A}}$ is an $L\times L$ matrix with each row containing only three non-zero elements corresponding to the prescription of the problem. As the possible outcomes can only be 0 or 11, the assigned variables can take the same values as well, $x_{i}=0,1$. The corresponding Hamiltonian is given by
$H_{P}=\sum_{c=1}^{M}\frac{1+J_{c}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}}{2},$ (40)
with alternating randomly as $J_{c}=\pm 1$ and with $M=L$. Following the strategy of Refs. [115, 13] we only consider such instances where the ground state is non-degenerate, allowing for a better characterization of the energy gap at the critical point. In these instances, one can drop the alternating coupling constant, $J_{c}=1$. The corresponding quantum annealing Hamiltonian is written as
@@
exhibiting an exponentially small gap around $\lambda_{c}=1/2$, where exponentially long driving times are required to reach near adiabatic dynamics. To this end, we probe the performance of both the sparsified QBCD Hamiltonian and the first-order locally optimized variational CD ansatz by the ground state fidelity and the minimal energy gap at the critical point. The first-order CD expansion is given by
@@
for which a similar system of equations can be derived based on the variational principle as in the case of the Max-Cut problem (see App. I).
Within the QBCD approach, we follow the same strategy similar to the previous subsection, Eq. (34) and determine the minimal gap, $\Delta$ and the corresponding eigenvectors, $\lvert\mathrm{GS}\rangle$ and $\lvert\mathrm{ES}\rangle$ with a mesh of solely $40$ points. Finally, the full QBCD was also sparsified to the same density as that of the leading order local CD expansion, Eq. (V.3).
We again use the ground-state gap and final fidelity to assess adiabaticity. As shown in Fig. 14, even for short annealing times ($T=5$), the local expansion amplifies the exponential gap decay only by a constant factor. In contrast, the sparsified QBCD achieves exponentially larger gaps with approximately polynomial decay, even for longer times ($T=10,20,80$), with its effect disappearing only once adiabaticity is reached in the bare process. Further details are shown about the relative gap amplification in App. G. Similar trends hold for the final fidelity (Fig. 15): at short times ($T=2$), QBCD significantly outperforms both the local expansion and the bare process, even when the latter runs much longer ($T=40$). In this regime, QBCD fidelities remain near unity, while the local expansion and bare dynamics still show exponential decay (see Fig. 14). Overall, the results for the XORSAT model further reinforce the conclusions drawn from the case of the Max Cut, highlighting both the efficiency and the prospects of the sparsified QBCD to boost currently realizable CD strategies.
Closing this subsection, we perform the analogous comparison of the local CD expansion and the sparsified QBCD for the XORSAT model in terms of energetic requirements. In this case, the Hilbert–Schmidt norm of the second-order CD term scales as
$\lvert\lvert H^{(2)}_{1}\rvert\rvert^{2}\sim L^{2},$ (43)
as shown in Fig. 16. The sparsified QBCD exhibits a similar trend to that observed for the Max Cut, with a slightly faster yet still subleading growth. This parallel energetic scaling further illustrates the Max Cut results, showing that sparsified QBCD significantly alleviates energetic demands for implementation
@@
Figure 16: Energy costs of the second and first order local CD expansions of the $3$-XORSAT and Max Cut problems quantified by the Hilbert-Schmidt norms growing approximately as $\sim L^{2}$, in agreement with Eq. (38) and Eq. (43). On the other hand, the sparsified QBCD exhibits a much smaller norm growing approximately linearly for the accessible system sizes.
@@ VI Discussion
In this study, we explored how quantum adiabatic dynamics can be enhanced to overcome spin-glass bottlenecks using various approximate counterdiabatic (CD) driving schemes. Specifically, we examined the variational local expansion and its improved version known as counterdiabatic optimized local driving (COLD). We also introduced a new scheme—quantum brachistochrone counterdiabatic driving (QBCD)—which focuses directly on the bottleneck region while relying only on the approximate knowledge of an exponentially small fraction of the full CD at a single parameter value. First, we tested the local expansion approaches on an exactly solvable minimal spin glass toy model described by an Ising chain admitting a free-fermionic description [99]. This model provides a valuable testbed with an exponentially small gap and two macroscopically distinct degenerate ground states. Its solvability gives access to larger system sizes and the extraction of asymptotic scalings, while the key ingredients and the level of efficiency observed here are further demonstrated in spin-glass models. To this end, we employed the variational principle to construct two- and three-body CD terms from the Floquet-Krylov expansion, which can efficiently be simulated on a classical computer [58, 59] and has also recently been realized on quantum hardware [109, 110, 120]. A key finding is that, while these CD ansatzes can exponentially amplify the size of the minimal gap, it remains exponentially small in the system size $L$, and the final ground state probability is of the same order of magnitude for times shorter than $\sim e^{\alpha L}$, with $\alpha$ given in Eq. (12).
@@
Our analysis highlights key features of CD of frustrated spin systems. First, from the gap characterization in Sec. IV.1, it can be inferred that CD can amplify the gap exponentially in the diabatic regime, and thus substantially reduce the diabatic transitions. However, reaching the regime of driving times where diabatic transitions come solely from the exponentially small ground state gap, the first two CD terms provide negligible improvement (see Fig. 6) and higher order expansions are required. On the other hand, slowing down the driving schedule also reduces the overall strength of the CD term, given the $\partial_{t}\lambda$ prefactor in Eq. (14). These features prevent the construction of an effective local CD term that ensures adiabatic dynamics in the presence of exponentially small gaps between frustrated energy levels.
In this context, CD techniques, including the variational formulation [58, 59, 60], aim to keep the evolved state close to the instantaneous ground state. In our simple model, a single transition spans the entire effective chain, Eq. (9), meaning the state must be adjusted non-locally to follow the changing ground state.
@@
As an alternative and novel approach, we introduced the QBCD Hamiltonian in Sec. V, which shares the structure of the Hamiltonian for the time-optimal evolution in the quantum brachistochrone problem for the corresponding localized edge states. The QBCD Hamiltonian employs only an exponentially small fraction of the full CD at a single parameter value, focusing exclusively on the ground and first excited states around the critical point. With the help of it, the success rate of the protocol increases from a vanishing probability to almost $0.5$, as indicated by the average final kink number. While implementing nonlocal, multi-body interactions on quantum hardware remains challenging, QBCD shows that even an exponentially small fraction of the full CD Hamiltonian can suffice for exponential speed-up: approximate knowledge at a single parameter point is enough, and the scheme remains robust near the critical point, drastically alleviating implementation complexity. The constant energetic cost of QBCD, in contrast to the diverging ones in the local expansions, further underscores its practical advantages, paving the way for experimentally feasible and resource-efficient CD expansion schemes in both quantum and classical simulations.
@@
In the second part of the paper, we examined how our findings extend to two realistic spin-glass models associated with NP-hard problems, namely the $3$-regular Max Cut and the $3$-XORSAT problems. These models exhibit an exponentially small gap already at system sizes $L\gtrsim 10$, accessible by classical numerical simulations. Specifically, the $3$-XORSAT problem extends the validity of our results to cubic spin-glass Hamiltonians. Due to the limitations in accessible system sizes, we used the final ground-state fidelity—rather than the excitation density—as a measure of adiabaticity, alongside the minimum gap. Our results show that local CD expansions have a similar efficiency as in the minimal spin-glass model. They yield only modest improvements outside the adiabatic regime: the gap is amplified primarily for fast quenches where, however, excitations to higher energy states break down adiabaticity completely. On the other hand, gap amplification terminates for slightly shorter driving time scales than those where the bare process fidelity starts to depart from zero. At the same time, for intermediate and slow annealing times, the fidelity under local CD driving preserves the same exponential decay with system size as in the bare dynamics, and thus no significant gains are obtained in this regime.
For the QBCD method, we introduced an enhanced, sparsified version that matches the density of the local expansions. Our findings show that exponential speed-up across an exponentially small gap can be achieved with only an exponentially small nonlocal contribution from the already substantially reduced full CD term at a single parameter value, with persistent efficiency under finite deviations from the critical point. In this sparsified scheme, only a polynomially-sized subset of eigenstate components needs to be retained, and the ground-state gap can be efficiently approximated using advanced large-scale numerical techniques such as quantum Monte Carlo methods [1, 2, 105] or reverse annealing [102, 103, 104], substantially lowering computational resources and facilitating large-scale classical implementations [117, 118]. From the viewpoint of quantum-device implementation, the controlled truncation of the residual non-locality guides the construction of approximate local CD terms with drastically reduced resource-requirements, enabling higher-order CD expansions on near-term quantum hardware [110, 109, 120].
@@
Finally, we also compared the efficiency of the COLD approach with the local expansion and the sparsified QBCD in the Max Cut problem. Although it provided a noticeable improvement over the local expansion, its performance remained consistently inferior to that of the sparsified QBCD. Even for small systems ($L=6$), spin-glass effects already reduce the efficiency of COLD below that of QBCD, and its numerical implementation proved challenging even at modest sizes ($L\lesssim 10$). As $L$ was increased and spin glass effects became stronger, the performance of COLD converged to that of the second-order local expansion. This highlights both its limited impact and the substantially higher computational costs compared to QBCD.
@@
This combination of results suggests that the challenge of traversing exponentially small gaps in these spin-glass problems is primarily of entropic rather than energetic origin [121]. In other words, overcoming the bottleneck does not require access to strong control fields but rather a minimal amount of information about where and how to steer the quantum state. This perspective further highlights the practical value of methods like QBCD, where even a small fraction of non-local terms is enough to achieve efficient evolution.
Our further results on the Hilbert–Schmidt norms in the spin-glass models demonstrate that sparsified QBCD is drastically less resource demanding than the corresponding local CD expansions. In both the Max Cut and XORSAT cases, the unfavorable polynomial growth of local CD expansions is replaced by a far more benign scaling under the sparsified QBCD. This improved behavior arises from the reduced non-locality and the sparsity of both $\partial_{\lambda}H$ and the eigenstates, which together suppress the divergences that make the norm of the exact QBCD diverge at critical points. Taken together, these results further support the notion that controlled truncation in sparsified QBCD can facilitate more resource-efficient implementations of approximate CD, providing a natural pathway to higher-order expansions.
@@
Accordingly, while exploring alternative strategies beyond a priori action minimization remains an interesting direction, our results—and the preceding discussion—suggest that concentrating on progressively improving approximate implementations of QBCD may ultimately be more effective for overcoming bottlenecks. Approaches based on the integration of CD ansatzes with optimal control methods—such as QAOA [63, 62, 64, 65]—optimize control fields iteratively using only the final state reached by the dynamics [77]. However, these methods often suffer from limited analytical transparency and can be hampered by complex control landscapes and barren plateaus [122, 123, 124, 125]. In contrast, the sparsified QBCD already demonstrates significant improvements and points towards a more efficient approximate implementation using experimentally accessible resources. Building on these advantages, future work could combine QBCD with complementary strategies, such as reverse annealing [102, 103, 104], bias fields [126, 87], Lyapunov control [67, 68], quantum Monte Carlo methods [1, 2, 105] or targeted control engineering near critical points.
@@ A.1 Jordan-Wigner transformation and Dirac fermions
$H[\lambda]=\sum_{j=1}^{L}\left[-\lambda J_{j}\tilde{\sigma}^{x}_{j}\tilde{\sigma}^{x}_{j+1}+(1-\lambda)\tilde{\sigma}_{j}^{z}\right].$ (46)
Tildes are used to distinguish the two sets of Pauli matrices for later convenience. This form of the Hamiltonian is amenable to a Jordan-Wigner transformation, mapping spins $\{\tilde{\sigma}_{j}\}$ to Dirac fermions $\{c_{j}^{\phantom{\dagger}},c_{j}^{\dagger}\}$ by means of the following identities,
@@
[…]
@@
[…]